whats the slope of a line that passes through the points (-3.5) and (7,10)
Answer by Maths68(1474)

Slope = rise in y / rise in x
slope = y2-y1 / x2-x1
Points (-3,5) and (7,10)
slope = (10-5)/(7-(-3))
slope = (10-5)/(7+3)
slope = 5/10
slope = 1/2